OpenStudy (haldko):

I'll fan and medal Find an equation for the nth term of the arithmetic sequence. a19 = -92, a20 = 6 Choices: an = -1856 + 98(n - 1) an = -1856 - 98(n - 1) an = -1856 - 98(n + 1) an = -1856 + 98(n + 1)

OpenStudy (haldko):

@sammixboo

OpenStudy (haldko):

@Hero @pooja195 @quickstudent PLEASE HELP

OpenStudy (welshfella):

to find the common difference between one term and the next we do the calculation a20 - a19 = 6-(-92) = ?

OpenStudy (sshayer):

find a1 by the formula \[a _{19}=a _{1}+(19-1)d\]